OpenStudy (anonymous):

In the basketball game, Sarah made three times as many 2-point shots as 3-point shots. She made 8 shots in all. How many 2-point shots did she make?

OpenStudy (anonymous):

8 shots in all and three times as many 2-pointers than 3-pointers, so she made six 2-pointer shots in total

OpenStudy (anonymous):

...

OpenStudy (anonymous):

let's see if your right

OpenStudy (anonymous):

okay :D

OpenStudy (anonymous):

sorry someone's bothering me right now XD

OpenStudy (anonymous):

it is correct because six 2-pointers and two 3-pointers makes eight shots in total, and you said she made 3 times as many 2-pointers than 3-pointers. 2x3=6. :D

OpenStudy (anonymous):

Hope i helped!

OpenStudy (anonymous):

ok back

OpenStudy (anonymous):

i reviewed the thingy and the answer was 6

OpenStudy (anonymous):

in otherwords correcto

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i read it wrong xD so got the question wrong

OpenStudy (anonymous):

O.o

OpenStudy (texaschic101):

x = 2 point shots y = 3 point shots x + y = 8 3y = x now we sub 3y + y = 8 4y = 8 y = 8/4 y = 2 3y = x 3(2) = x 6 = x she made 6 two point shots and 2 three point shots